Power of point source is 18mW. Find the intensity 2.5 m away from source.

Grade:12th pass

1 Answers

Susmita
5 years ago
Intensity is power transferred per area.Again area is proportional to length square.
Intensity=P/r2=18×10-3/(2.5)2=2.88 mW/m2
if it heps please approve.